OpenStudy (anonymous):

Find the empirical formula: 0.31g of phosphorus reacts to form 0.71g of phosphorus oxide. The oxide has a relative molecular mass of 284.

OpenStudy (abmon98):

4P+5O2-->P4O10 Use the equation Number of moles=Mass(g)/Molar Mass(g/mol) 0.31/30.9=0.0100 0.71/284=0.0025 Divide by the smallest figure 0.0100/0.0025=4 moles of Phosphurus 0.0100/0.0100=1 moles of P4O10